# Time Transfer: On Optimal Learning Rate and Batch Size In The Infinite Data Limit

### v1

@@ 4 Discussion
While originally, we were aiming to find a golden recipe for hyperparameter transfer in the infinite data limit, we show that there is no simple and straight-forward answer. However, we do believe that there exists a deeper underlying perspective on the problem, as opposed to the one of simply tuning learning rate and batch size.
@@
Fundamentally, the field of model parametrization research has originated from and is further converging towards preserving some notion of norm in some infinite (model width and/or depth) limit (Everett et al., 2024; Yang et al., 2024; Large et al., 2024). In fact, any parametrization itself is simply a set of scaling rules to be applied to hyperparameters in order to preserve these norms (e.g. of model weight matrices or weight updates). Expanding on this, one can argue that scaling rules follow from the requirement of keeping some underlying quantity invariant within the infinite limit. From this perspective, hyperparameter transfer is nothing but a consequence of such “conservation laws”.
@@
With this perspective in mind, we draw a parallel between infinite model and data limits, and speculate that a similar notion of “norm” should exist and should be aimed to be preserved in the infinite data limit. In fact, there is already a good candidate for this, namely the noise scale (Eq. 4 and 8), which intriguingly also induces scaling rules for hyperparameters (see Appendix A.2 for in-depth discussion). However, the existing definition neither takes into account the adaptive nature of the optimizer, nor the scenario of jointly following the infinite data and model limits. From our experimental observations, the following definition might be an applicable extension:
$B_{\mathrm{noise}}^{\mathrm{Adam}}\propto\eta_{\mathrm{eff}}\frac{\sqrt{T}}{\sqrt{\frac{B_{\mathrm{crit}}(T,M)}{B}}+\sqrt{\frac{B}{B_{\mathrm{crit}}(T,M)}}},$ (1)
where $B_{\mathrm{crit}}(T,M)$ is the critical batch size as a function of time $T$ and model size contribution $M$ via the corresponding model parametrization, as motivated by Park et al. (2019); $\eta_{\mathrm{eff}}=\eta(\eta_{\mathrm{max}},\beta_{1},\beta_{2})$ is the effective learning rate, incorporating additional dependency on Adam’s $\beta_{1,2}$ parameters. In our experiments with $\mu$P, it appears that $B_{\mathrm{crit}}(T,M)=B_{\mathrm{crit}}(T)$ as we do not observe any dynamics changing across the $\mu$P model family (Sec. 3.3). Once the invariant is established, one can derive the corresponding scaling rules for $T,M,B,\eta$ in the joint $(M,T)\to\infty$, effectively resulting in the hyperparameter transfer.
@@
The main limitation of our work is that we provide only experimental hints for Eq. 1. This formula is the fruit of empirical observation only, which makes it a scaling law and should not be taken as a theorem. Nonetheless, we hope that our results make the first step towards the unification of infinite data and model size limits via deriving such a joint scaling invariant, also inclusively across multiple model parametrizations. Finally, our insights into optimal scaling rules for learning rate and batch size might be valuable for practitioners who approach the problem of hyperparameter optimization in the infinite data and model size limit. We provide our summary and recommendations in Appendix A.1.
@@ 5 Related Work
$\boldsymbol{(\eta,B)}$ scaling rules    In efforts to accelerate model training, the $\eta\propto B$ rule for the SGD optimizer was found necessary to avoid performance loss due to increased batch size (Goyal et al., 2018), known as generalization gap (Keskar et al., 2017). Afterwards, additional usage of momentum (Smith et al., 2018) and model scaling (Park et al., 2019) was incorporated, and a $\eta\propto\sqrt{B}$ rule for Adam was observed (Hilton et al., 2022). From the theoretical side, experimentally observed rules were verified with the framework of stochastic differential equations (SDEs) (Smith & Le, 2018; Malladi et al., 2023), loss curvature analysis (Zhang et al., 2019; McCandlish et al., 2018; Li et al., 2024) and random matrix theory (Granziol et al., 2021). While most of the studies were performed in the fixed epoch budget, Shallue et al. (2019) broadened the perspective to other target budget measures and studied the scope of the $\eta\propto B$ rule applicability across various datasets and model architectures. Looking beyond fixed budgets, Smith & Le (2018) showed a linear relation between the optimal batch size and the dataset size (for fixed $\eta$), and Smith et al. (2020) similarly presented hints for a linear relation between the optimal learning rate and the dataset size (for fixed $B$), with both works considering the SGD optimizer. In the modern LLM pretraining context, Hu et al. (2024); DeepSeek-AI et al. (2024) approached this problem by deriving the joint $(\eta,B)$ scaling laws.
$\boldsymbol{\mu}$P    Originally developed within the Tensor Program series studying feature learning in the infinite width limit (Yang & Hu, 2022; Yang et al., 2022), $\mu$P has been gaining traction recently within the LLM community. It has been extensively tested and applied experimentally (Lingle, 2024; Blake et al., 2024; Gunter et al., 2024; Dey et al., 2024), as well as theoretically, with Yang et al. (2023); Bordelon et al. (2024) extending it to the infinite depth limit, and Yang et al. (2024); Bernstein et al. (2023) revisiting it from the spectral normalization perspective. Recently, Everett et al. (2024) showed that other model parametrizations also induce hyperparameter transfer if taking weight alignment into account. Furthermore, they revealed that $\mu$Transfer does not work in the regime of Chinchilla-optimal scaling (Hoffmann et al., 2022). The most closely related work to ours, Shen et al. (2024) expanded on this observation and proposed a learning rate scheduler combining $\mu$P and experimentally measured $(\eta,B)$ scaling rules to allow for the hyperparameter transfer in the $T\to\infty$ limit, however only limited to the $\eta^{*}\propto 1/\sqrt{T}$ scaling regime.
@@ 6 Conclusion
In this work, we studied joint model and data scaling in the LLM context from the perspective of optimal learning rate and batch size dynamics. We observed an intricate interplay of three optimal learning rate scaling regimes in the infinite data limit, controlled by the batch size in its relation to the critical batch size as it evolves in time. This dynamic is preserved during model scaling with $\mu$P, as well as the loss sensitivity to the learning rate variation, highlighting the intriguing difference in how $\mu$P infinite width and time limits evolve the critical batch size. Overall, we hope our observations pave the way towards deeper understanding of the optimal scaling in the unified infinite data and model size limit.
@@ A.1 Hyperparameter optimization in the infinite data and model size limit
    After the grid search, one should be able to obtain a plot similar to Fig. 1(b) and Fig. 2(a). Provided long enough WSD horizon, a drift in time of the critical batch size region, associated to the peak of the fixed token budget curve in Fig. 1(b), should be visible. Likewise, there should be a drift of the optimally tuned (i.e. assuming optimal learning rate is used) batch size in time as in Fig. 2(a). Since we observe a strong correlation but still a mismatch between the optimally-tuned batch size and the critical batch size, we suggest the following approach for selecting optimal hyperparameter values:
@@
      Extrapolate the drift of the fixed budget curve in time based on Fig. 1(b) (in our case, approximately $B^{\mathrm{crit}}\propto T$) and derive the expected critical batch size for the target horizon $B^{\mathrm{crit}}_{\mathrm{target}}$.
@@
      Set optimal learning rate for the target horizon as:
      $\eta^{*}_{\mathrm{target}}=\begin{cases}\vspace{2mm}\eta^{\mathrm{crit}}\cdot\sqrt{B^{*}_{\mathrm{target}}/B^{\mathrm{crit}}_{\mathrm{target}}}&\text{if $B^{*}_{\mathrm{target}}\leq B^{\mathrm{crit}}_{\mathrm{target}}$}\\ \vspace{2mm}\eta^{\mathrm{crit}}\cdot\sqrt{B^{\mathrm{crit}}_{\mathrm{target}}/B^{*}_{\mathrm{target}}}&\text{if $B^{*}_{\mathrm{target}}>B^{\mathrm{crit}}_{\mathrm{target}}$}\end{cases},$ (2)
      where we correct the learning rate value for the corresponding scaling regime, assuming the optimum learning rate corresponding to the critical batch size remains constant in time $\eta^{\mathrm{crit}}=\eta^{*}(T)|_{B^{\mathrm{crit}}(T)}=\mathrm{const}$.
@@
We suppose it is also possible to adjust the recipe above to the continual learning setting (Çağatay Yıldız et al., 2024; Ibrahim et al., 2024): under assumption of $\eta^{\mathrm{crit}}$ being constant in time and of the golden path hypothesis (Vyas et al., 2024), one could indefinitely run the model training with the same learning rate but dynamically adjust the batch size to follow the critical one (peak of the fixed budget curve in Fig. 1(b)), or, alternatively viewed, to remain on the pareto curve of Fig. 2(b) (inset plot).
@@ A.2 On critical batch size and noise scale
where $T$ is the training set size. It is suggested that one should aim at finding the optimal noise scale in the first place, rather than optimal batch size and learning rate. Within the suggested Bayesian framework, Smith & Le (2018) argue that the optimality arises from the trade-off between depth and breadth in the Bayesian evidence. In a follow-up work, Park et al. (2019) take one step further and extend the noise scale to a model width limit and introduce a modified noise scale accounting for the change of the model width in the standard (SP) and Neural Tangent Kernel (NTK) parametrizations (Jacot et al., 2020):
@@
  • GeLU activation function, Layer Normalization initialized with $1$ (Ba et al., 2016), RoPE with $\theta=10000$ (Su et al., 2023).
@@
A.5 Random seed variation
@@
A.6 Learning rate schedule scaling (cont.)
@@
A.7 Loss profiles per $(d_{\mathrm{model}}^{\mathrm{base}},d_{\mathrm{model}})$ configuration
A.7.1 $d_{\mathrm{model}}^{\mathrm{base}}=1024,~{}d_{\mathrm{model}}=1024$
@@
A.7.2 $d_{\mathrm{model}}^{\mathrm{base}}=1024,~{}d_{\mathrm{model}}=512$
@@
A.7.3 $d_{\mathrm{model}}^{\mathrm{base}}=1024,~{}d_{\mathrm{model}}=256$

### v2

@@ 4 Discussion
While originally, we were aiming to find a golden recipe for hyperparameter transfer in the infinite data limit, we show that there is no simple and straight-forward answer. In Sec. 3.2, we show that the model fit based on Eq. 1 describes the observed data points well. However, as illustrated in Sec. 3.3, following the optimal learning rate trajectory in time is not sufficient to obtain optimal performance. That leads us to believe that there exists a deeper underlying perspective on the problem, as opposed to the one of simply tuning learning rate and batch size.
@@
Fundamentally, the field of model parametrization research has originated from and is further converging towards preserving some notion of norm in some infinite (model width and/or depth) limit (Everett et al., 2024; Yang et al., 2023a; Large et al., 2024). In fact, any parametrization itself is simply a set of scaling rules to be applied to hyperparameters in order to preserve these norms (e.g. of model weight matrices or weight updates). Expanding on this, one can argue that scaling rules follow from the requirement of keeping some underlying quantity invariant within the infinite limit. From this perspective, hyperparameter transfer is nothing but a consequence of such “conservation laws”.
@@
With this perspective in mind, we draw a parallel between infinite model and data limits, and speculate that a similar notion of “norm” should exist and should be aimed to be preserved in the infinite data limit. In fact, there is already a good candidate for this, namely the noise scale (Eq. 5 and 9), which intriguingly also induces scaling rules for hyperparameters (see Appendix A.2 for in-depth discussion). However, the existing definition neither takes into account the adaptive nature of the optimizer, nor the scenario of jointly following the infinite data and model limits.
@@
We hope that our experimental observations, similarly to the discovery of the $\eta^{*}\propto T$ scaling rule for SGD, will make the first step towards the theoretical unification of infinite data and model size limits via deriving such a joint scaling invariant. Still, our insights into optimal scaling rules for learning rate and batch size might be valuable for practitioners who approach the problem of hyperparameter optimization in the infinite data and model size limit. We provide our recommendations in Appendix A.1.
As future work, it would important to improve the resolution in data points with a finer grid of $(\eta,B)$ values. This is a necessary step to establish the generalization power of Eq. 1 and the power law fits of its critical parameters as a function of time (Fig. 2), also across various data sets, model architectures and modalities. Additionally, while we used $\mu P$ as the main way to incorporate model scaling due to its ability to transfer optimal learning rate across model sizes, recent work of Everett et al. (2024) suggests that this is not the only way to do so. A similar study to ours, but for other model parametrizations, is an exciting direction of future research.
@@ 5 Related Work
$\boldsymbol{(\eta,B)}$ scaling rules    In efforts to accelerate model training, the $\eta\propto B$ rule for the SGD optimizer was found necessary to avoid performance loss due to increased batch size (Goyal, 2017), known as generalization gap (Keskar et al., 2016). Afterwards, additional usage of momentum (Smith, 2017) and model scaling (Park et al., 2019) was incorporated, and a $\eta\propto\sqrt{B}$ rule for Adam was observed (Hilton et al., 2022). From the theoretical side, experimentally observed rules were verified with the framework of stochastic differential equations (SDEs) (Smith & Le, 2017; Malladi et al., 2022), loss curvature analysis (Zhang et al., 2019; McCandlish et al., 2018; Li et al., 2024) and random matrix theory (Granziol et al., 2022). While most of the studies were performed in the fixed epoch budget, Shallue et al. (2019) broadened the perspective to other target budget measures and studied the scope of the $\eta\propto B$ rule applicability across various datasets and model architectures. Looking beyond fixed budgets, Smith & Le (2017) showed a linear relation between the optimal batch size and the dataset size (for fixed $\eta$), and Smith et al. (2020) similarly presented hints for a linear relation between the optimal learning rate and the dataset size (for fixed $B$), with both works considering the SGD optimizer. In the modern LLM pretraining context, Hu et al. (2024); Bi et al. (2024) approached this problem by deriving the joint $(\eta,B)$ scaling laws.
$\boldsymbol{\mu}$P    Originally developed within the Tensor Program series studying feature learning in the infinite width limit (Yang & Hu, 2020; Yang et al., 2022), $\mu$P has been gaining traction recently within the LLM community. It has been extensively tested and applied experimentally (Lingle, 2024; Blake et al., 2024; Gunter et al., 2024; Dey et al., 2024), as well as theoretically, with Yang et al. (2023b); Bordelon et al. (2023) extending it to the infinite depth limit, and Yang et al. (2023a); Bernstein et al. (2023) revisiting it from the spectral normalization perspective. Recently, Everett et al. (2024) showed that other model parametrizations also induce hyperparameter transfer if taking weight alignment into account. Furthermore, they revealed that $\mu$Transfer does not work in the regime of Chinchilla-optimal scaling (Hoffmann et al., 2022). The most closely related work to ours, Shen et al. (2024) expanded on this observation and proposed a learning rate scheduler combining $\mu$P and experimentally measured $(\eta,B)$ scaling rules to allow for the hyperparameter transfer in the $T\to\infty$ limit, however only limited to the $\eta^{*}\propto 1/\sqrt{T}$ scaling regime.
@@ 6 Conclusion
In this work, we studied joint model and data scaling in the LLM context from the perspective of optimal learning rate $\eta$ and batch size $B$ dynamics. We observed an intricate dependence of optimal $\eta$ scaling on $B$ and its relation to the critical batch size $B_{\mathrm{crit}}$, as a function of the pretraining token budget $T$. This dynamic is preserved during model scaling with $\mu$P, as well as the loss sensitivity to the learning rate variation, highlighting the intriguing difference in how $\mu$P infinite width and time limits evolve the critical batch size. Overall, we hope our observations pave the way towards deeper understanding of the optimal scaling in the unified infinite data and model size limit.
@@ A.1 Hyperparameter optimization in the infinite data and model size limit
    After the grid search, one should be able to obtain a plot similar to Fig. 1(a) and Fig. 3(a). Provided long enough WSD horizon, a drift in time of the critical batch size region, associated to the peak of the fixed token budget curve in Fig. 1(a), should be visible. Likewise, there should be a drift of the optimally tuned (i.e. assuming optimal learning rate is used) batch size in time as in Fig. 3(a). Since we observe a strong correlation but still a mismatch between the optimally-tuned batch size and the critical batch size, we suggest the following approach for selecting optimal hyperparameter values:
@@
      Perform a fit to fixed-budget curves per token budget step based on data similar to Fig. 1(a) with Eq. 1, following the procedure of Sec. 3.1. Fit a set of extracted $B_{\mathrm{crit}}$ per token budget with a power law function $B_{\mathrm{crit}}=aT^{\alpha_{B}}+b$ to extract the corresponding exponent $\alpha_{B}$ (in our case, $\alpha_{B}\approx 1$) and derive the expected critical batch size for the target horizon $B^{\mathrm{crit}}_{\mathrm{target}}$.
@@
      Perform the same power law fit to the $\eta^{\mathrm{crit}}(T)$ data and extrapolate its value to the target horizon, obtaining $\eta^{\mathrm{crit}}_{\mathrm{target}}$. Set optimal learning rate for the target horizon as:
      $\eta^{*}_{\mathrm{target}}=\begin{cases}\vspace{2mm}\eta^{\mathrm{crit}}_{\mathrm{target}}\cdot\sqrt{B^{*}_{\mathrm{target}}/B^{\mathrm{crit}}_{\mathrm{target}}}&\text{if $B^{*}_{\mathrm{target}}\leq B^{\mathrm{crit}}_{\mathrm{target}}$}\\ \vspace{2mm}\eta^{\mathrm{crit}}_{\mathrm{target}}\cdot\sqrt{B^{\mathrm{crit}}_{\mathrm{target}}/B^{*}_{\mathrm{target}}}&\text{if $B^{*}_{\mathrm{target}}>B^{\mathrm{crit}}_{\mathrm{target}}$}\end{cases},$ (3)
      where we correct the learning rate value for the corresponding $\eta^{*}(B)$ scaling regime.
@@
We suppose it is also possible to adjust the recipe above to the continual learning setting (Yıldız et al., 2024; Ibrahim et al., 2024): under assumption of $\eta^{\mathrm{crit}}$ being constant in time and of the golden path hypothesis (Vyas et al., 2023), one could indefinitely run the model training with the same learning rate but dynamically adjust the batch size to follow the critical one (peak of the fixed budget curve in Fig. 1(a)), or, alternatively viewed, to remain on the pareto curve of Fig. 3(b) (inset plot).
@@ A.2 On critical batch size and noise scale
where $T$ is the training set size. It is suggested that one should aim at finding the optimal noise scale in the first place, rather than optimal batch size and learning rate. Within the suggested Bayesian framework, Smith & Le (2017) argue that the optimality arises from the trade-off between depth and breadth in the Bayesian evidence. In a follow-up work, Park et al. (2019) take one step further and extend the noise scale to a model width limit and introduce a modified noise scale accounting for the change of the model width in the standard (SP) and Neural Tangent Kernel (NTK) parametrizations (Jacot et al., 2018):
@@
  • GeLU activation function, Layer Normalization initialized with $1$ (Lei Ba et al., 2016), RoPE with $\theta=10000$ (Su et al., 2024).
@@
A.5 Fitting procedure
In Sec. 3.1, we introduce the fitting procedure to the data points of Fig. 1(a) with a functional form of Eq. 1 with two parameters $\eta_{\mathrm{crit}}$ and $B_{\mathrm{crit}}$. We perform the fit separately per token budget, as illustrated in Fig. 6, using scipy.optimize.curve_fit() and firstly without including error bars (which stem from the variation of the model with $\mu P$ and random seeds). Since we observe some data points as having no uncertainties, which makes the fit computationally unstable, we repeat the fit two more times: one with adding small $\epsilon=10^{-15}$ as uncertainty for such data points. Then, we attribute the mean uncertainty across the other points to the points without uncertainties and perform the same fit. This procedure results in three data sets for each $\eta_{\mathrm{crit}}(T)$ and $B_{\mathrm{crit}}(T)$, corresponding to three variations of the fitting procedure, which we treat as “systematic” uncertainty. For each data point, we assign an uncertainty as a square root of the corresponding covariance matrix element, as obtained from the fit.
We then perform a full power law $p_{\mathrm{crit}}=a_{p}T^{\alpha_{p}}+b_{p},~{}p\in\{\eta,B\}$ fit with uncertainties to each of the three data sets, for each of the critical parameters $p$. We obtain the following values with corresponding fit uncertainties:
  • No error:
    • $(a_{\eta},\alpha_{\eta},b_{\eta})=(1.9\cdot 10^{5},-0.85,2.9\cdot 10^{-3}]$
    • $(a_{B},\alpha_{B},b_{B})=[8.2\cdot 10^{-5},1.00,3.0\cdot 10^{5})$
    • $(\sigma_{a_{\eta}},\sigma_{\alpha_{\eta}},\sigma_{b_{\eta}})=(1.2\cdot 10^{6},0.32,2.8\cdot 10^{-4})$
    • $(\sigma_{a_{B}},\sigma_{\alpha_{B}},\sigma_{b_{B}})=(4.5\cdot 10^{-4},0.23,1.5\cdot 10^{5})$
  • With errors + $\epsilon$:
    • $(a_{\eta},\alpha_{\eta},b_{\eta})=(2.3\cdot 10^{9},-1.31,2.7\cdot 10^{-3})$
    • $(a_{B},\alpha_{B},b_{B})=(2.3\cdot 10^{-2},0.75,1.9\cdot 10^{5})$
    • $(\sigma_{a_{\eta}},\sigma_{\alpha_{\eta}},\sigma_{b_{\eta}})=(1.0\cdot 10^{10},0.21,1.1\cdot 10^{-4})$
    • $(\sigma_{a_{B}},\sigma_{\alpha_{B}},\sigma_{b_{B}})=(5.1\cdot 10^{-2},0.09,8.8\cdot 10^{4})$
  • With errors + mean uncertainty attribution:
    • $(a_{\eta},\alpha_{\eta},b_{\eta})=(4.5\cdot 10^{12},-1.68,2.9\cdot 10^{-3})$
    • $(a_{B},\alpha_{B},b_{B})=(4.9\cdot 10^{-7},1.20,2.8\cdot 10^{5})$
    • $(\sigma_{a_{\eta}},\sigma_{\alpha_{\eta}},\sigma_{b_{\eta}})=(4.5\cdot 10^{13},0.47,1.4\cdot 10^{-4})$
    • $(\sigma_{a_{B}},\sigma_{\alpha_{B}},\sigma_{b_{B}})=(2.1\cdot 10^{-6},0.17,8.0\cdot 10^{4})$
Since we are primarily interested in the power exponents $\alpha_{p}$, we average their values across the three variations to produce a central value for each $p$. For the uncertainties, we add in quadratures the variance across the three fit variations and the mean uncertainty obtained from each of the individual fits. This produces the results we supply in the main text (Eq. 2). To further constrain large uncertainties on $a_{p}$, we refit $\eta_{\mathrm{crit}}(T)$ and $B_{\mathrm{crit}}(T)$ time dependence data with the power exponents $\alpha_{p}$ fixed to the ones obtained above. This gives us the final model parameters, which we visualize oi Fig. 2 and 1(b) and discuss throughout the main text:
  • $a_{\eta}=(2.0\pm 0.3)\cdot 10^{9}$
  • $\alpha_{\eta}=-1.3\pm 0.4$
  • $b_{\eta}=(3.1\pm 0.1)\cdot 10^{-3}$
  • $a_{B}=(8.0\pm 1.3)\cdot 10^{-5}$
  • $\alpha_{B}=1.0\pm 0.2$
  • $b_{B}=(3.0\pm 1.1)\cdot 10^{5}$
Figure 6: Fits to token budgets $T={2^{30},2^{31},\ldots,2^{37}}$ (from upper left to bottom right) with Eq. 1 to the data points in Fig. A.9.
A.6 Random seed variation
@@
A.7 Learning rate schedule scaling (cont.)
@@
A.8 Loss profiles per $(d_{\mathrm{model}}^{\mathrm{base}},d_{\mathrm{model}})$ configuration
A.8.1 $d_{\mathrm{model}}^{\mathrm{base}}=1024,~{}d_{\mathrm{model}}=1024$
@@
A.8.2 $d_{\mathrm{model}}^{\mathrm{base}}=1024,~{}d_{\mathrm{model}}=512$
@@
A.8.3 $d_{\mathrm{model}}^{\mathrm{base}}=1024,~{}d_{\mathrm{model}}=256$